OpenStudy (anonymous):

Find the x-intercept and y-intercept. 2x+y=7

OpenStudy (anonymous):

hey.. replace x by 0 to get y intercept replace y by 0 to get x intercept

OpenStudy (anonymous):

for example replacing x by 0 gives: 2*0+y = 7 y =7 it's the y intercept

OpenStudy (anonymous):

I know that,But I'm confused after that. It's like 2(0)+y=7 Right? But I don't know what to do after that...

OpenStudy (anonymous):

Oh..Okay..I get it..I think..

OpenStudy (anonymous):

But what about the x? Is it 2x+0=7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x=7 what is x then ?

OpenStudy (anonymous):

5? If I'm supposed to subtract 2 from both sides...

OpenStudy (anonymous):

I hate Pre-Algebra>.<

OpenStudy (anonymous):

no divide both sides by 2 x = 7/2 x=3.5

OpenStudy (anonymous):

How do you know when to divide or subtract?..

OpenStudy (anonymous):

2x = 7 it means 2*x = 7 so you must divide in order to get rid of the multiplication but when there is addition you have to subtract

OpenStudy (anonymous):

Oh..Okay..

OpenStudy (anonymous):

2+x=7 here you have to subtract 2 from both sides

OpenStudy (anonymous):

And then if there's subtraction you add right?

OpenStudy (anonymous):

yes! x-2=7 add 2 to both sides

OpenStudy (anonymous):

if there is a division you have to multiply x/2 = 7 multiply by 2

OpenStudy (anonymous):

I get it..I have another thing..It says Write this equation in slope-intercept form, and then find the slope and y-intercept. -y=3x So what do I do? I mean,I know the answer is -3,) But I don't know how to find it.

OpenStudy (anonymous):

divide both sides by -1 and get y = -3x in such equations (y=ax+b) a means slope and b means y intercept

OpenStudy (anonymous):

So where does the -1 come from?